40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2694     Entry time: Fri Nov 12 14:21:32 2021
Author: Stephen, Radhika 
Type: DailyProgress 
Category: Cryo vacuum chamber 
Subject: Upgrade to Rigid Copper Bar, and assorted transitions from PD testing 

This post describes upgrade efforts from 10 - 16 November, with the following goals:

 - introducing a solid copper bar thermal linkage

 - shifting the setup away from PD testing

 - preparing for the next test (radiative cooling of Si)

Here are some highlights of the effort:

  • While removing the shields we found contaminants had plated near the line of sight surfaces at the optical window and the electrical feedthrough [Attachment 1]. The film was removed by IPA wipe [Attachment 2] and was not evident in any other location (presumably these were the cold surfaces in line of sight, so they received the most contaminant, but there may be a thinner deposition throughout!).
  • In order to install the copper rod, we needed to cut out slots in the outer shield and inner shield. We used a reciprocating saw and held the piece stiffly on a table [Attachment 3] [Attachment 4].
    • We tried to use large snips, but that failed to provide enough cutting force, especially where it was necessary to use the tip to access into the flanged corner. We also damaged a few of the electrical leads to the feedthrough (formerly used to wire the OSEMs) - this will require attention at next opportunity.
  • The copper rod itself was found to have a few issues:
    • Outer surfaces were somewhat tarnished and greasy, so an 80 grit aluminum oxide paper was used to clean up all surfaces [Attachment 5]. The surfaces were then wiped with IPA using Alpha wipes.
    • Slot width was matching the long thermal strap instead of the short thermal strap, so a drill press was used to add cut away the correct areas of one pair of holes [Attachment 6]. Later, this modification required larger washers in the stack under the bolt head.
    • The cold head had a larger OD than designed, so we were unable to use the corner holes to mount the RTD. There was not enough space at the corners to host the nut or washer. Instead one of the bolt pattern holes was used to host the RTD stack [Attachment 7]
  • The installation of the copper rod required the following steps:
    • We documented the previous state of the table [Attachment 8].
    • We removed the inner shield, the outer shield, and the old thermal linkage.
      • To access the cold head side of the linkage, we had to remove the lower conflat of the T.
    • We installed a 6" conflat flange, double faced with 4" bore (Lesker p/n DFF600X400), between the cryocooler and the T. This spacer raised the bottom face of the cold head to the correct height, so that the rigid copper bar runs approximately through the center of the shield apertures.
      • This required an order of new bolts with 3" length, to squeeze the three flanges (cryocooler, spacer, T) that are now stacked together.
    • We bolted the rigid copper bar to the coldhead, yawing the cryocooler to match the conflat bolt hole orientation while also pointing the copper bar down the axis of the arm.
      • This used new vented screws (UC Components p/n C-412-A) which are also silver plated, except in the location borrowed for the RTD stack as mentioned above [Attachment 7]. The RTD stack included a nut to adjust spring compression independently from screw threading.
      • Apiezon N thermal grease was applied on both surfaces to improve thermal conductivity across this joint.
      • We initially forgot to reinstall the mylar sheet radiation shielding that had been removed from the area around the cold head and around the linkage. This required that we reopen the bottom conflat to install the coldhead mitten, and that we pitch the aluminum shields away from the rigid bar to allow the mylar sheet to be inserted from the inside.
      • We found that the coldhead RTD had failed during intial mounting efforts, and a new RTD (actually one that had been desoldered from the setup previously, removed from the workpiece on 2021.08.07 but found to have no issue) was soldered and attached to the cold head, under the spring clamp.
    • Each shield was reinstalled, including:
      • The newly cut slots were used to pass shields over the rigid copper bar.
      • Electrical cabling was threaded through the usual apertures.
      • The outer shield was positioned on the G10 spacers.
      • Rough alignment was completed based on clearance from the rigid copper bar, and line of sight to optical window.
    • Final touchups were implemented:
      • Aluminum foil covered unused outer shield apertures.
      • A small aluminum foil panel was placed underneath the rigid copper bar, to cover the slot in the inner shield.
      • Final clamping of the inner shield and the heater (with indium gasket) were completed.
      • Thermal strap was used to link the cold baseplate to the rigid copper bar, with a bolted joint at the copper bar and a dog clamp joint at the baseplate. Apiezon N grease was applied to all contact faces.

The rest of the installation effort is captured in the next log post QIL/2695, to partition the items relevant to the radiative cooling of the silicon mass.

The photos here (and others) are posted to the QIL Cryo Vacuum Chamber photo album.

Attachment 1: IMG_0352.JPG  2.651 MB  | Hide | Hide all | Show all
IMG_0352.JPG
Attachment 2: IMG_0353.JPG  2.675 MB  | Hide | Hide all | Show all
IMG_0353.JPG
Attachment 3: IMG_0371.JPG  2.691 MB  Uploaded Mon Nov 15 08:33:48 2021  | Hide | Hide all | Show all
IMG_0371.JPG
Attachment 4: IMG_0370.JPG  2.631 MB  Uploaded Mon Nov 15 08:34:21 2021  | Hide | Hide all | Show all
IMG_0370.JPG
Attachment 5: IMG_0343.JPG  2.727 MB  Uploaded Mon Nov 15 08:40:51 2021  | Hide | Hide all | Show all
IMG_0343.JPG
Attachment 6: IMG_0351.JPG  2.937 MB  Uploaded Mon Nov 15 08:42:55 2021  | Hide | Hide all | Show all
IMG_0351.JPG
Attachment 7: IMG_0358.JPG  1.990 MB  Uploaded Mon Nov 15 09:01:28 2021  | Hide | Hide all | Show all
IMG_0358.JPG
Attachment 8: IMG_0346.JPG  2.767 MB  Uploaded Mon Nov 15 09:01:49 2021  | Show | Hide all | Show all
Attachment 9: IMG_2734.jpeg  164 kB  Uploaded Tue Nov 16 16:39:53 2021  | Hide | Hide all | Show all
IMG_2734.jpeg
Attachment 10: IMG_2740.jpeg  190 kB  Uploaded Tue Nov 16 16:40:04 2021  | Hide | Hide all | Show all
IMG_2740.jpeg
ELOG V3.1.3-